One coin is randomly selected from a jar containing 70 nickels, 100 dimes, 80 quarters, and 50 one-dollar coins. Find each probability p(quarter or nickel)
Found 2 solutions by math_tutor2020, greenestamps:
Answer by math_tutor2020(3817)


70+100+80+50 = 300 coins total

Of those 300 coins, 70+80 = 150 are either a quarter or a nickel.

150/300 = 1/2 is the final answer as a fraction.

Answer by greenestamps(13200)



P = (# of good outcomes) / (total number of outcomes)

P(quarter or nickel) = (number of quarters and nickels) / (number of coins) = (150)/(300) = 1/2

ANSWER: 1/2

Convert to decimal or percent if needed